# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v3

@@ 1.4 Hypotheses of Function-Driven Discretisation
Additionally, it is essential to note that isotropic networks still enable such clustering. They retain the ability through their bias terms, but they are suggested not to disproportionately encourage it through task-agnostic biases; instead, it would likely reflect a task-necessitated collapse. Hence, there is no reason to expect isotropy in representational distributions, as the underlying data or latent layer may include discretised aspects that arise independently of model effects. Therefore, this is not a test of whether distributions are anisotropic or isotropic; instead, it is the investigation of whether the emergent structure specifically about a distinguished direction can be turned on or off through functional form choices. Overall, isotropic networks may be expected to produce emergent structures that are more representative of the underlying data, rather than being dependent on functional forms.
Other speculative pathological modes are attributed to the hypothesised functional form quantisation. For example, discrete representations may be particularly susceptible to perturbations which move activations outside of the cluster, into an in-between region of representation space. Maps from these unpopulated regions may be poorly trained and fragile, due to their training-time rarity, and therefore highly unpredictable in their map. This may suggest that such a quantising inductive bias decreases adversarial robustness Goodfellow et al. (2015); Moosavi-Dezfooli et al. (2017); Ilyas et al. (2019); Tsipras et al. (2019), whereas removing this bias through isotropy may be advantageous. However, this speculation is not explored within this work.
This paper aims to assess the validity of the first prediction. The motivating questions are: Do discrete-symmetry defined functional forms tend to produce approximately discrete representations? This would be a quantising of an otherwise underlying continuous data structure. This is contrasted with continuous-symmetry defined functional forms, which are expected to tend to preserve a more continuous representation. Hence, this would demonstrate that such choices can carry representational inductive biases, with isotropy’s frame independence being expected to be a more preferable and natural choice, without arbitrarily distinguished directions. Overall, determining the validity of this prediction may provide crucial insight into the potential of Isotropic deep learning and the emergence of downstream interpretability structure, namely alignment, grandmother neurons, general interpretable directions, discrete codes, and possibly Superposition Elhage et al. (2022) and perhaps some connection to Neural Collapse Papyan et al. (2020).
A secondary outcome is an indication of whether the discrete construction can be pathological in nature. Since quantised representations may harm expressiveness, isotropic functions may help in tasks where this expressiveness must be retained. Preliminary evidence resulting from the above primary investigation suggests that function-driven representational collapse can be pathological in nature, leading to increased task error. However, this study is not intended to elucidate the exact mode of the pathology, such as those proposed above, so the results remain a preliminary indication.
Consequently, this study can be considered to be determining a more foundational prerequisite phenomenon than the Spotlight Resonance Method observed and causally explained. This prior method elucidated how functional forms can causally influence alignment structure. In contrast, this work investigates whether forms can lead to the emergence of any induced task-agnostic structure. Determining whether current forms induce representational quantisation may indicate a causal mechanism for these other downstream phenomena.
It is established that discretisation can be contingent on the functional form’s maximal algebraic symmetries. It is shown that these symmetries are good predictors of the emergent task-agnostic structure in representations. Overall, the establishment of these links suggests these functional forms are not benign in their effect and can induce unintended biases into networks indicative of their respective maximal symmetries. It would encourage the identification and documentation of the extent of these influences, since leveraging such properties could be advantageous. It suggests that considering maximal symmetries at the primitive scale may be important, and motivates a reevaluation of the contemporary functional form of primitives in general.
@@ 3 Experiments and Interpretation
In this section, comparisons between isotropic and anisotropic activation functions are displayed and discussed. The examples shown are an excerpt from a larger variety of results available in App. B. The exact method for training these networks is discussed in App. D. A code repository for the PPP method is available at https://github.com/GeorgeBird1/Spotlight-Resonance-Method.
@@
Figure 1: Displays ten rows of PPP-method’s results, divided into columns. Each row represents an independent autoencoder network trained on the reconstruction of the CIFAR dataset Alex (2009). Each column represents the results of the PPP method at various stages of training. The leftmost column shows a freshly initialised model before training, and moving rightward, the network is progressively trained for up to 125 epochs, as shown in the rightmost column. Hot-spots indicate where the internal latent layer is particularly dense with representations — collated over all samples from the CIFAR training set. The top five rows depict networks which utilise the anisotropic activation function, standard-tanh, whilst the bottom five rows utilise the isotropic activation function, isotropic-tanh. Every other detail is identical otherwise. Figure titles indicate the exact number of epochs trained for. This specific network consists of a latent layer of 18 neurons, with standard (unnormalised) input-output pairs drawn from CIFAR. The dark centres about the origin are attributed to a vanishing volume due to an angular threshold, rather than the absence of representations.
@@
Firstly, the construction of this specific autoencoder ensured no activation function preceded the latent layer of study, similarly to the spotlight resonance method. This eliminated the possibility that these alignment phenomena are trivially due to the bounding geometry of an activation function, including limit points. The activation functions are only applied after the latent layer of study. Hence, this indicates how the functional forms in later layers can exert a representational bias generally through optimisation throughout the network. Thus, the functional form can bias the parameter trajectories during learning, which then transforms the representations into these discrete structures. Therefore, this represents a non-trivial representational bias due to functional forms, confirming the prior prediction Bird (2025a), and in the general case preceding any trivial bounding or neural refraction. Further results also substantiate this.
@@
Additionally, the results suggest that discrete-representational Superposition Elhage et al. (2022) may arise uniquely in the anisotropic functional form. Several rows, namely two through five, display discrete directions that are either aligned with or anti-aligned with the standard basis, but are sometimes also arranged at more general angles. This is further demonstrated in Fig. 2, which is the same autoencoder but with a greater number of hidden latent layers.
@@
This is believed to be indicative of a discrete representational structure consistent with descriptions of Superposed features [^4] These are thought to be representational examples of the phenomenon, which differs from the original measurement, which was often concerning the autoencoding network’s parameters. Thus, this potentially provides a sort of dual observation of the phenomenon. Additionally, row two demonstrates exclusively anti-aligned representations to the standard basis, which appears to diverge from the classical alignment-first features. Elhage et al. (2022). Additionally, it also shows that activation functions can exert influence backwards through the model, affecting representations in latent layers that precede any activation function in the forward pass. This reframes several Superposition experiments Elhage et al. (2022), which identify the phenomenon in such cases.
@@
This may indicate that discrete Superposition phenomena arise due to anisotropic forms, but whether these forms provide a complete description remains debatable — especially for the phenomenon’s dual through parameters, which could still be applicable under continuous representations. Further analysis of individual planes from the PPP method may be necessary to assess the nuances of the Superposition in more detail, rather than relying on this conglomerate plot across all planes. However, this would require many more samples to populate individual planes sufficiently. Additionally, several columns on isotropic networks for early-stage training also indicate directions of high density, but these progressively distribute throughout training. These are believed to be statistical in nature, as optimisation has not had the opportunity to transfer asymmetry in forms to representations yet. This is discussed further in App. A. Overall, it is suggestive that representational Superposition may be a network-adaptive emergent phenomenon predicated upon the imposed task-agnostic anisotropic biases.
@@
Generally, anisotropic functional forms appear to produce an increasingly discrete structure in representations as training progresses, starting with angularly arranged but more diffuse representations and gradually refining into narrow, angular beams. This appears consistent with a linear feature hypothesis Elhage et al. (2022). This structure can be seen to be contingent on the anisotropic functional form, suggesting it carries a quantising inductive bias. Additionally, the anisotropic network’s linear features align about both positive and negative distinguished directions, appearing to respect the sign-flip and permutation (hyperoctahedral symmetry) of the standard-tanh functional form. This may be indicative of the analytically degenerate orthants produced by such forms. This is contrasted with the isotropic form examples, which retain a general continuous representational structure which becomes decreasingly discrete through training.
@@
A final observation is that spontaneous symmetry breaking appears to set distinct directions early during training, and later training only subtly refines these directions. This indicates a strong dependence on initialisation in networks and supports considering the initialiser’s probabilistic symmetry presented in the symmetry-based taxonomy Bird (2025a). These may also constitute inductive biases through interactions. Additionally, in the future, this could be explored in terms of whether isotropic networks have a slightly reduced dependence on initialisation, due to momentum enabling a traversal of identically flat connected basins along the symmetry group’s orbit Bird (2025a).
@@
Figure 2: Displays an identical plot to Fig. 1, except for the results being drawn from a network with three latent layers, each with 18 neurons per layer. The latent layer studied is the first latent layer, which precedes any activation function.
@@
The appendices present a multitude of additional results for autoencoders of varying widths, depths, datasets, input-output normalisations, latent layer and activation-function dependencies. All of these indicate a similar general tendency towards task-agnostic structure in representations arising from the choice of functional form, particularly the algebraic symmetry involved. Moreover, they also suggest that geometries in representational superposition may be more likely a result of symmetry than the orthogonal to antipodal to Thomson geometries, which are suggested to balance representation capacity and interference Elhage et al. (2022). This is because hyperoctahedral tanh forms continue to demonstrate orthogonal arrangements first, yet Leaky-ReLU permutation-only forms move immediately to antipodal arrangements for the same task. This would encourage a reconsideration of function-derived symmetries as the primary causal mode for the particular geometries produced, when balancing interference and representational capacity.

### v4

@@ 1.3 Discussion of Methodological Approach
It is essential to note that isotropic networks still enable such clustering. They retain the ability through their bias terms, but they are suggested not to disproportionately encourage it through task-agnostic biases. Instead, it would likely reflect a task-necessitated collapse. Hence, there is no reason to expect perfect isotropy in representational distributions, as the underlying data or latent layer may include discretised aspects that arise independently of model effects. This makes the symmetry-induced structure the pivotal signature of these function-induced effects.
Therefore, this is not a test of whether distributions are anisotropic or isotropic; instead, it is the investigation of whether the emergent structure, specifically about distinguished directions, can be turned on or off through functional form choices. Overall, isotropic networks may be expected to produce emergent structures that are more representative of the underlying data, rather than being dependent on functional forms. This frame independence may be generally preferable and a natural choice, without task-agnostic distortions induced by distinguished directions. Hence, determining the validity of this prediction may provide crucial insight into the potential of Isotropic deep learning as an approach.
This work also advances the prior Spotlight Resonance Method (SRM). This modified and novel tool is called the “Privileged-Plane Projective Method” (PPP method) and operates on a similar principle.
However, unlike SRM, this new and adapted tool does not project out the vector length in the process. Thus, information regarding both angular and magnitude distributions is retained. This provides a more nuanced understanding of the network’s complicated representational structures, which is critical for better interpreting the organised structure and the phenomena of study.
Additionally, the PPP method does not produce any angular-falloff smearing effect related to the cone width of SRM. Therefore, this modified method enables higher-fidelity angular maps of representations, which may be beneficial in many circumstances and essential for determining the exact nature of representational structures [^3] Together, the magnitude-preservation and high-fidelity angular discernment, uniquely position the Privileged-Plane Projective method as a tool for assessing whether neural refraction [4] occurs for out-of-training-distribution outlier representations.. Together, this high-resolution geometric and intuitive insight, which importantly preserves magnitude-dependent, is the methodological intent behind developing and introducing the PPP visualisation tool used in this study.
This also positions the two tools — the existing SRM and the newly introduced PPP — for this study as similar, complementary, yet distinct approaches. The former remains better suited to statistical measures of angular representations, whilst the latter is preferable where complex geometry and high-resolution discernment are critical — such as in this study. The loss of angular fidelity in angular clustering metrics, such as SRM, can lead to a loss of discernment between the discrete linear features and the superposed arrangements of interest in this study. Hence, the novel PPP was developed and prioritised in this study to determine such structures, revealing a clearer geometric structure of representations. The following Sec. 3 provides a detailed overview of this new tool.
Using a comparative analysis between standard-Tanh and isotropic-Tanh, and similarly for the Leaky-ReLU family of functions, this PPP method can determine whether functional-form symmetry constitutes a causal mechanism for a tendency for discrete representation cluster structure to form due to task-agnostic functional-form symmetries.
It is used to demonstrate that algebraic permutation equivariance tends to produce a symmetry-organised quantising inductive bias on otherwise, underlying, near-continuous data structure, rather than task-necessitated representational collapses. This is contrasted with continuous-symmetry-defined functional forms, which are expected to tend to preserve a more continuous representation by not introducing spurious structure. Validation would directly demonstrate that such choices can carry representational inductive biases and require significant exploration.
Consequently, it suggests that downstream phenomena, such as interpretable neurons, discrete codes, (representational) superposition [8], and possibly Neural Collapse [21], are not fundamental but rather a direct consequence of anisotropic functional forms. This may have benefits for AI safety and interpretable networks; however, preliminary evidence also suggests that this function-driven, task-agnostic collapse and structure can be pathological, leading to increased reconstruction error.
These pathological qualities of anisotropy are a secondary, preliminary outcome of this work. It may be associated with reduced representational capacity and expressiveness, as suggested by results in App. B, and with other hypothesised causes, such as robustness, discussed in App. C.3. However, this study is not intended to elucidate the exact mode of the pathology, so the results remain a preliminary indication. As a future direction, this may be investigated, resulting in a performance-interpretability tradeoff, if findings continue to be supported.
2 Theoretical Framework and Considerations
In this section, the theoretical aspects of the predictions are significantly developed. Particularly articulating how and why detectable signatures may form, their predicted appearance, alongside their potential implications.
A further discussion of how the theory may relate to other architectures and of the potential pathological implications of functional-form biases is presented in App. C, which provides further theoretical details.
2.1 Mechanistic Discussion of Function-Induced Structure
In this section, a more detailed discussion of how the proposed phenomenon is hypothesised to occur in networks is provided. This discussion makes clear the identifiable and detectable signatures that may present as a consequence of the theory. These explain why the prediction of quantised representations was expected for anisotropy, and lack thereof for isotropy, deriving from previous predictions [4].
A further discussion pertaining to gradient-flow and first-order representational corrections is provided in App. C.2, which examines the mathematical emergence of the phenomenon.
This section’s discussion is felt to deepen and ultimately supersede the more simplistic ‘distinguished direction’ heuristic. This is because, multivariately, the notion of distinguished directions is rather absolute, spatially discrete, and arbitrary in how they are defined. Instead, one can consider the multivariate maps to be more spatially continuous and locally varying in the regions they favour and to what degree they are preferred. Hence, this generalises the finite distinguished directions to a more spatially-continuous ‘priviledging-field’-like approach. Where the priviledging-field is stronger, representations accumulate; where lower, they disperse — although the staticness of such a field is not implied. Notably, distinguished directions do remain important in characterising the particular representation of the (discrete) symmetries, so are retained to define the planes of the PPP method.
At a high level, the inducement of discrete structure is argued to stem from the function’s anisotropy with its uneven computational map. Differing localities of this map may have differing practical usabilities. The network may act to leverage these by moving representations towards favourable areas and away from mappings of less utility — or, in the most general sense, attractors and repulsers may naturally form that influence representations’ trajectories during optimisation. Hence, one would generally expect anisotropy to have more and less probabilistically favourable regions for representations to be embedded by the parameterised maps — where absolute uniformity is possible but probably unlikely to form.
Consequently, one would expect the anisotropy to shift from the algebraic definition, in the functional form, into the representations to some extent. These would yield predictable disturbances in density as the network learns to leverage the more useful local transforms. The implication is that these representations may gather more densely into certain regions whilst rarifying others during optimisation. These would be function-induced structures that form predictable over- and under-represented areas. These become a baseline signature of the phenomenon for that function.
Note that this does not determine the number of such clusters, which depends on the particular function that instantiates the form, nor, at this stage, its organised structures or its generalisability beyond the specific function.
This motivates the second essential aspect: the symmetry of this functional form. This determines the nature of that anisotropy, particularly its organisation, through its computational map’s degeneracies. Additionally, symmetry is also important not only in the inducement of such organised structure and implication for representational degrees-of-freedom, but this organisation also relates to the functional form as a whole. It may allow collation of generalisable behaviours across forms in a taxonomised symmetry approach [4]. The determination of such generalised findings would be important for model design along a symmetry-primitive design axis. Generalised findings supporting this taxonomy are demonstrated within this paper. A novel avenue could also explore general functions that can then be decomposed into a unique selection of symmetry-defined functional forms with known biases, each of which might contribute to the combined function.
Beginning with the symmetry organisation. Depending on the symmetry, these favourable computational maps may exhibit discrete degeneracy — locally repeating at numerous instances throughout the representation space. This may be beneficial to the network whilst providing a clear and detectable signature.
If a particular map is highly advantageous, then the network may be expected to crowd many representations across its localised non-linearity to achieve its desired computation. If there is only one occurrence of such a map, then the network may balance this desirable computation whilst retaining sufficient representational degrees of freedom. If it becomes too concentrated, information encoded by the distances between representations may be forfeited or suppressed as clusters become increasingly concentrated.
For example, this may take the form of many directions being projected onto a single axis, reducing the distribution’s dimensionality and leading to conflation of associated semantics. Another possibility is that specific directions tangential to the usable map are scaled down to bring them closer to the usable locality, suppressing the associated semantics. These examples are non-exhaustive heuristics. This may be similarly applicable symmetry-or-not444‘Not’ still relates to the identity symmetry, you just lose the crucial signature of symmetry-dictated organisation. Similarly, in the SRM study [5], some proposed maps don’t clearly arise from symmetry for some (approximate) Thomson bases, yet they are likely to still have some discrete rotational symmetry, even if approximate; hence, representation organisation may still occur about them. Potentially different discrete symmetries arise at different approximation tolerances. This may be explored through weighted combinations of functional forms..
Hence, the network may redistribute representations to balance global representational degrees of freedom that express varied semantic information as a tradeoff with the locality of usable maps. Some nuance remains: some fraction of representations may benefit from one form of map, whilst another set may benefit from a different locality, potentially forming multiple clusters.
However, with the symmetry of such functional forms, there is this symmetry-induced organised degeneracy to these usable maps. The balance between representational degrees of freedom and densifying over these desirable maps is a markedly different tradeoff. The network needn’t sacrifice inter-activation structure to the same degree; instead, important representational degrees of freedom can be retained by exploiting these same/similar maps in differing localities, such that the distribution remains more distinctly separable and more semantics are preserved.
The consequence of this hypothesis is that the denser patches of activations develop a task-agnostic structure, as clusters recur over similar maps, which are themselves organised directly by this symmetry of the functional form. The result is recurrent, symmetry-organised densification of representations onto multiple equivalent patches, without sacrificing degrees of freedom as much as concentrating in just one place. This organised densification is the key signature arising from such symmetry, which the PPP method then amplifies into observables.
The construction of the ablation trial can validate this empirically. One would expect to see a symmetry-defined angular structure in anisotropic networks, since the practicality of local maps fluctuates; however, in isotropic networks, one would not expect to see such task-agnostic, form-driven structure, as all angular localities are equivalent — there is no need for the network to overrepresent certain areas. It is the continuous angular limit in map degeneracy; there is no need to suppress angular degrees of freedom when seeking a localised map.
Hence, the tradeoff between representational degrees-of-freedom and leveraging localised computations disappears in isotropic networks, as there are no angularly localised maps [^5] This makes clear the goal of isotropic function search: to recreate a highly usable map across all angles and hence such a tradeoff is never necessitated. Anisotropic maps necessarily vary because the maximal symmetry is discrete, so this is not possible; the network must strike a balance. This positions isotropy as potentially preferable if suitable functions can be found. If multiple maps are needed for differing representations, one can stack isotropic functions and potentially explore function-routing masks per representation/locality. — it can utilise practical mappings without a detrimental tradeoff [^6] A radial tradeoff may still occur, due to the non-linear map along any radial direction. However, some non-linearity is required for operation. This associates with this paper’s proposed coding scheme: a magnitude-direction hypothesis for stimulus-degree and semanticity, respectively.. Hence, suppression of particular semantics needn’t happen unless task-necessitated.
Consequently, one would predict that the general signature of discrete-symmetry networks is a task-agnostic representational structure organised about the defining symmetry. A distinct absence of such task-agnostic structure would then be expected for isotropy. Observing a difference in task-agnostic structure between these would validate this general hypothesis. This general approach will be used to study how various function-driven symmetries interact and act on the network and its representations, determining the various biases induced.
For example, considering permutation-like symmetries about a particular basis would result in an orthant partitioning of the space [4]. Depending on the specific permutation symmetry, various sets of these orthants may be analytically degenerate, producing transformed copies of equivalent maps — a manifestation of the underlying group definition. The network’s optimisation may then shape representations about these orthants, producing analogous clusters under these equal mappings and yielding an approximate symmetry structure resembling the orthant decomposition and the initial, defining permutation symmetry. Hence, a response of optimisation to the symmetry.
Generalising these anisotropic structures may be an important foundational mode for the formation of classical axis-alignment, discrete linear features, (representational) Superposition, grandmother neurons and more.
It is also singularly essential to note that this is a form-driven ‘task-agnostic’ organisational structure, something seen repeated through different networks and hence infers a foundational function-driven bias as opposed to a data-dependent fluctuation. Data may naturally contain such over- and underdensities, so the detection of which is not sufficient to deem it a foundational bias. Hence, conflation between the algebraic an/isotropy of the function and a direct expectation of a distributional an/isotropy in representations is not encouraged.
Instead, the tendency towards repeated, and symmetry-organised, structure is the hallmark signature of this potentially significant and unappreciated bias. This motivates the careful setup of this study to demonstrate clearly that it is function-driven, not data-driven, by detecting repeated instances of the same structure within the same data and task, then comparing them against differing functional forms, which constitutes the comparative ablation approach.
The PPP method is particularly effective in amplifying this symmetry-organised structure by conglomerating many single planes, themselves defined by the symmetry, into a composite plot. Hence, if such over- and under-densities are organised by symmetry, then their discrepancies will compound, making the signature more evident.
This does not eliminate the possibility of a single, very strong, dense patch of representations, rather than many smaller symmetry-organised patches. This latter consideration constitutes an additional reason why a comparative ablation trial against isotropy is essential, as either form of task-agnostic structure can be attributed to the functional form presence of anisotropy if it does not similarly occur within the purely isotropic networks. Despite this caveat, in this paper, symmetry remains the central defining factor shown to induce structure. It will be shown that PPP frequently demonstrates several clearly visible patches organised about the symmetry. Hence, the results advance the symmetry-based functional form theory on how the wide variety of interpretable structures fundamentally form in a causal manner. This makes clear these are consequences of symmetry-defined functional form choices than any fundamental and spontaneous emergent/mysterious occurrences of such representational phenomena.
Finally, a note on the sharpness of such clusters. In general, these may vary between diffuse and sharp clusters, depending on the map, but should remain symmetry-organised fluctuations in any case due to the shared functional form.
In the extreme of a highly usable locality near much less usable localities, one would expect to observe much sharper representation distributions, producing near-discrete-like quantised clusters. The shape and narrowness of these may indicate the shape and extent of the useful computations. This may produce narrow beams where inter-activation degrees of freedom are suppressed in favour of the improved map, alongside the semantics associated with these directions, as separability is lost.
The number of such beams is still undetermined by the symmetry-defined form, as is specific to the function, but may in future work be explored in terms of a multiplicity due to the symmetry (particularly in the upcoming orthant picture). In the results, it is observed that multiple narrow, discrete-like beams shift and converge into a single stronger beam, showing that optimisation forms highly dense patches over particularly useful maps, which may shift to continue ‘finding’ better transforms.
It also so happens, that the observations provided in this work, do tend to have these very sharp nearly-discrete quantised representational clusters for anisotropy, and often approximatly smooth continua for isotropy, making the analysis incidently more idealistic and hence, more certainty in conclusions — although, both do show some variation from plot-to-plot. However, despite an association between the two, it may not always be expected that anisotropic and isotropic definitions for functions and representation distributions will coincide.
Overall, it is instead the symmetry-led organisation of generally more concentrated and rarefied representation patches which is the general signature. These structured unevennesses in the representations would then reflect the function-form-induced symmetry. Where multiple functions and symmetries are involved, this structure may become more complicated; hence the necessity of a minimalistic, otherwise isotropically controlled autoencoder on reconstruction to minimise these complicating confounders.
This section offers a mechanistic account predicting that functional forms are not benign in their effects on representations, but instead directly induce their structure into predictable, task-agnostic fluctuations in the density of representations. The organisation of these directly stems from the symmetry defining the particular functional form. Drawn from this is the primary hypothesis that since isotropic functional forms are devoid of distinguished directions, they may consequently eliminate any tendency towards task-agnostic representational collapse organised by a discrete symmetry’s uneven map.
Validating this unequivocally shows that foundational biases arise directly from our choices of functional form for primitives, and other primitives must be similarly investigated. Since these choices are then shown to carry consequences, this necessitates a highly significant interrogative analysis of the foundational functions of the deep learning approach and supports a taxonomisation of functions and their consequences. This generalised investigation is encouraged to span across initialisation, normalisers, optimisers, operations and more. If its findings generalise, they change the conceptual foundation for how we understand the organisation and appearance of learned representations. This study supports the hypothesis that such choices may be significant representational inductive biases and worth considerable exploration.
2.2 Broader Impact Concerns
Although indirect, there exists potential for the consequences of isotropy to be leveraged for undesirable means. As it will be shown, isotropic networks tend to produce smoother continua of representations, opposed to anisotropic’s inducement of clustered representational structure. With isotropic development comes the risk of intentionally using such functions to create influential networks that are functionally black boxes. This is a critical issue where networks may need to be transparent in decision-making, such as in the medical or justice systems.
Many current interpretability tools rely on this cluster separability to intuit model decisions, predicted behaviours, and broadly understanding the network — isotropic functions induce distributions opposed to such tooling. They could be used to obfuscate such understanding, maliciously making the models harder to audit, and in turn requiring the development of new tooling to match existing interpretability.
It is felt that practitioners should be mindful of these non-negligible risks when implementing functions and, in parallel, strive to develop tooling that brings the interpretability of isotropic networks in line with current anisotropic tooling. This may also be circularly advantageous for creating flexible, generalised tools that may benefit the broader field beyond isotropy. This ensures that developments in isotropic model capability do not come at the expense of transparency.
Both the SRM [5] and the PPP method already provide examples of isotropic-considerate interpretability tools. Incorporating built-in semantic interpretation into these tools may give further insight for safety-critical applications.
Overall, the development of isotropic functions is encouraged as a practical and advantageous route to exploring and leveraging foundational biases; yet, it is vital to be cautious of the stated downsides and proactively mitigate this potential eventuality.
@@ 4 Experiments and Interpretation
In this section, comparisons between isotropic and anisotropic activation functions are displayed and discussed. The examples shown are an excerpt from a larger variety of results available in App. B. The exact method for training these networks is discussed in App. E. A code repository for the PPP method is available at https://github.com/GeorgeBird1/Spotlight-Resonance-Method.
Additionally, a small disclaimer: in the extended results of App. B, a normalisation technique is used to mitigate dataset-specific statistical artefacts arising from the $\left[0,1\right]$ bounding of data elementwise. This normalisation step consists of a preprocessing step in which each sample has the dataset pixel-wise mean subtracted and the pixel-wise variance unit-standardised, excluding pixels that are zero. This reduced the impact of data-format-related artefacts on the findings. This process is discussed more thoroughly in App. A.
However, this preprocessing standardisation was considered somewhat artificial and was introduced for pragmatic reasons. Therefore, it is not included in the results of this section. However, the observations were present with and without normalisations. Thus, conclusions reached are not influenced by such choices, as quantisation induced by anisotropic functions was seen in both cases, with only minor qualitative differences.
@@
Figure 1: Displays ten rows of PPP-method’s results, divided into columns. Each row represents an independent autoencoder network trained on the reconstruction of the CIFAR dataset [1]. Each column represents the results of the PPP method at various stages of training. The leftmost column shows a freshly initialised model before training, and moving rightward, the network is progressively trained for up to 125 epochs, as shown in the rightmost column. Hot-spots indicate where the internal latent layer is particularly dense with representations — collated over all samples from the CIFAR training set. The top five rows depict networks which utilise the anisotropic activation function, standard-tanh, whilst the bottom five rows utilise the isotropic activation function, isotropic-tanh. Every other detail is identical otherwise. Figure titles indicate the exact number of epochs trained for. This specific network consists of a latent layer of 18 neurons, with standard (unnormalised) input-output pairs drawn from CIFAR. The dark centres about the origin are attributed to a vanishing volume due to an angular threshold, rather than the absence of representations.
@@
Firstly, the construction of this specific autoencoder ensured no activation function preceded the latent layer of study, similarly to the spotlight resonance method. This eliminated the possibility that these alignment phenomena are trivially due to the bounding geometry of an activation function, including limit points. The activation functions are only applied after the latent layer of study. Hence, this indicates how the functional forms in later layers can exert a representational bias generally through optimisation throughout the network. Thus, the functional form can bias the parameter trajectories during learning, which then transforms the representations into these discrete structures. Therefore, this represents a non-trivial representational bias due to functional forms, confirming the prior prediction [4], and in the general case preceding any trivial bounding or neural refraction. Further results also substantiate this.
@@
Additionally, the results suggest that discrete-representational Superposition [8] may arise uniquely in the anisotropic functional form. Several rows, namely two through five, display discrete directions that are either aligned with or anti-aligned with the standard basis, but are sometimes also arranged at more general angles. This is further demonstrated in Fig. 2, which is the same autoencoder but with a greater number of hidden latent layers.
@@
This is believed to be indicative of a discrete representational structure consistent with descriptions of Superposed features [^8] These are thought to be representational examples of the phenomenon, which differs from the original measurement, which was often concerning the autoencoding network’s parameters. Thus, this potentially provides a sort of dual observation of the phenomenon. Additionally, row two demonstrates exclusively anti-aligned representations to the standard basis, which appears to diverge from the classical alignment-first features. [8]. Additionally, it also shows that activation functions can exert influence backwards through the model, affecting representations in latent layers that precede any activation function in the forward pass. This reframes several Superposition experiments [8], which identify the phenomenon in such cases.
@@
This indicates that discrete Superposition phenomena arise due to anisotropic forms, but whether these forms provide a complete description remains debatable — especially for the phenomenon’s dual through parameters, which could still be applicable under continuous representations. Further analysis of individual planes from the PPP method may be necessary to assess the nuances of the Superposition in more detail, rather than relying on this conglomerate plot across all planes. However, this would require many more samples to populate individual planes sufficiently. Additionally, several columns on isotropic networks for early-stage training also indicate directions of high density, but these progressively distribute throughout training. These are believed to be statistical in nature, as optimisation has not had the opportunity to transfer asymmetry in forms to representations yet. This is discussed further in App. A. Overall, it is suggestive that representational Superposition may be a network-adaptive emergent phenomenon predicated upon the imposed task-agnostic anisotropic biases.
@@
Generally, anisotropic functional forms appear to produce an increasingly discrete structure in representations as training progresses, starting with angularly arranged but more diffuse representations and gradually refining into narrow, angular beams. This appears consistent with a linear feature hypothesis [8]. This structure can be seen to be contingent on the anisotropic functional form, suggesting it carries a quantising inductive bias. Additionally, the anisotropic network’s linear features align about both positive and negative distinguished directions, appearing to respect the sign-flip and permutation (hyperoctahedral symmetry) of the standard-tanh functional form. This may be indicative of the analytically degenerate orthants produced by such forms. This is contrasted with the isotropic form examples, which retain a general continuous representational structure which becomes decreasingly discrete through training.
@@
A final observation is that spontaneous symmetry breaking appears to set distinct directions early during training, and later training only subtly refines these directions. This indicates a strong dependence on initialisation in networks and supports considering the initialiser’s probabilistic symmetry presented in the symmetry-based taxonomy [4]. These may also constitute inductive biases through interactions. Additionally, in the future, this could be explored in terms of whether isotropic networks have a slightly reduced dependence on initialisation, due to momentum enabling a traversal of identically flat connected basins along the symmetry group’s orbit [4].
@@
Figure 2: Displays an identical plot to Fig. 1, except for the results being drawn from a network with three latent layers, each with 18 neurons per layer. The latent layer studied is the first latent layer, which precedes any activation function. Moreover, one can see in rows 1, 3, 4, and 5 that narrow individual beams slowly shift in position during optimisation, converging on strong axis-aligned distributions. This may suggest these axis-aligned positions offer more favourable non-linear maps for computation than elsewhere.
@@
The appendices present a multitude of additional results for autoencoders of varying widths, depths, datasets, input-output normalisations, latent layer and activation-function dependencies. All of these indicate a similar general tendency towards task-agnostic structure in representations arising from the choice of functional form, particularly the algebraic symmetry involved. Moreover, they also suggest that geometries in representational superposition may be more likely a result of symmetry than the orthogonal to antipodal to Thomson geometries, which are suggested to balance representation capacity and interference [8]. This is because hyperoctahedral tanh forms continue to demonstrate orthogonal arrangements first, yet Leaky-ReLU permutation-only forms move immediately to antipodal arrangements for the same task. This would encourage a reconsideration of function-derived symmetries as the primary causal mode for the particular geometries produced, when balancing interference and representational capacity.
It is also established that discretisation can be contingent on the functional form’s maximal algebraic symmetries. Across all results, these maximal symmetries are shown to be good predictors of the emergent task-agnostic structure in representations. The establishment of these links suggests these functional forms are not benign in their effect and can induce significant, unintended biases into networks indicative of their respective maximal symmetries. This initiates and further encourages the identification and documentation of these influences and their extent.
These results show that considering maximal symmetries at the primitive scale may be important, and motivate a reevaluation of the contemporary functional form of primitives in general. These result observations collectively reframe the foundational origin of discrete structure in representations and the assumptions therein. Leveraging such properties could be highly advantageous.